### v3

@@ Abstract
Advances in transformer-based language models have highlighted the benefits of language-specific pre-training on high-quality corpora. In this context, German NLP stands to gain from updated architectures and modern datasets tailored to the linguistic characteristics of the German language. GeistBERT seeks to improve German language processing by incrementally training on a diverse corpus and optimizing model performance across various NLP tasks. It was pre-trained using fairseq with standard hyperparameters, initialized from GottBERT weights, and trained on a large-scale German corpus using Whole Word Masking (WWM). Based on the pre-trained model, we derived extended-input variants using Nyströmformer and Longformer architectures with support for sequences up to 8k tokens. While these long-context models were not evaluated on dedicated long-context benchmarks, they are included in our release. We assessed all models on NER (CoNLL 2003, GermEval 2014) and text classification (GermEval 2018 fine/coarse, 10kGNAD) using $F_{1}$ score and accuracy. The GeistBERT models achieved strong performance, leading all tasks among the base models and setting a new state-of-the-art (SOTA). Notably, the base models outperformed larger models in several tasks. To support the German NLP research community, we are releasing GeistBERT under the MIT license.
keywords:
Language Modeling, Machine Learning for NLP, Resources and Evaluation
@@ 1 Introduction
The advancement of neural language modeling (LM) in natural language processing (NLP) has been driven by the development of contextual pre-trained word representations, particularly through transformer-based architectures. Models like Bidirectional Encoder Representations from Transformers (BERT) [9] have significantly impacted the field by providing robust, generalized representations that can be fine-tuned for specific downstream tasks, enhancing performance across various NLP applications. While much of the early work focused on English and multilingual models, it has become clear that single-language models, particularly those trained on large, high-quality corpora, can outperform their multilingual counterparts when applied to their target language.
Building on this understanding, the German NLP community has seen the introduction of models like GottBERT [26], which leveraged the German portion of the OSCAR [21] corpus to create a high-performance RoBERTa-based [18] model tailored specifically for the German language. However, as the field evolves, so too must the approaches to model training. Recent developments in pre-training methodologies, such as Whole Word Masking (WWM) [7] and the availability of newer, more extensive corpora like OSCAR23 [15], OPUS [28], and mC4 [34], present opportunities to further refine and enhance German language models.
RoBERTa-based models like GottBERT are often limited to a 512-token input length, which can be insufficient for long documents and broader contextual understanding. To address this, we extend the input length to 8k tokens using Longformer [2] and Nyströmformer [32], which mitigate the computational constraints of full self-attention. We adapted the models using roberta2longformer [^1] https://github.com/LennartKeller/roberta2longformer. Longformer replaces full self-attention with local and task-specific global attention, reducing complexity while preserving contextual awareness. It remains compatible with RoBERTa, allowing seamless weight transfer from pre-trained models [18]. Nyströmformer approximates self-attention using the Nyström method for low-rank matrix decomposition, eliminating the need for specialized attention patterns. This improves scalability and flexibility, without the need for additional pre-training, making it a strong alternative for handling longer sequences.
To fully leverage these architectural improvements for German NLP, we introduce GeistBERT, a German Enhanced Incremental Semantic Tuning BERT model. GeistBERT builds on the foundation laid by the best checkpoints of filtered GottBERT models (i.e. ${}^{\mathrm{f}}\mathrm{GottBERT}$), incrementally pre-training on top of it using a combination of partly deduplicated modern German corpora, including OSCAR23, mC4 from CulturaX [20], Wikipedia and corpora from OPUS. By integrating WWM and harnessing the diverse, expansive nature of these contemporary corpora, GeistBERT aims to set a new benchmark for German language models, offering improved performance and broader applicability across a range of NLP tasks.
@@
  • We incrementally trained GeistBERT on top of ${}^{\mathrm{f}}\mathrm{GottBERT}$ using a combination of partly deduplicated modern German corpora (OSCAR23, OPUS, mC4), OpenLegal and Wikipedia.
  • We integrated Whole Word Masking (WWM) into the pre-training process to enhance the model’s ability to capture semantic relationships within the German language.
  • We provide GeistBERT as a base model as well as Longformer and Nyströmformer variants to the community, accessible under an open-source license for further usage.
GeistBERT represents a significant step forward in the development of German-specific transformer models, offering enhanced capabilities through modern training techniques and expansive, high-varying data. By extending the input length to 8k tokens through Longformer and Nyströmformer, our model is well-suited for handling long-text tasks, enabling more effective document-level language understanding and processing.
@@ 2 Related Work
The rise of transformer-based models like BERT [9] marked a major shift in NLP, enabling significant performance improvements. Originally introduced as an English model and later as a multilingual version (mBERT), BERT’s success led to monolingual adaptations tailored to specific languages. For German, models like GermanBERT [^2] https://www.deepset.ai/german-bert and dbmdz BERT [^3] https://huggingface.co/dbmdz/bert-base-german-uncased emerged, trained on datasets of 12GB–16GB, sourced from Wikipedia, news articles, and legal texts.
RoBERTa enhanced BERT by training on a larger 160GB corpus, optimizing the architecture, and removing next sentence prediction. This strategy was applied to other languages, resulting in models like CamemBERT [19] for French and RobBERT [8] for Dutch, highlighting the benefits of large, diverse training corpora.
In German NLP, GBERT and GELECTRA [5] built on this progress by training on 145GB of the OSCAR corpus [21] and additional sources, surpassing earlier German BERT models. These advancements underscored the impact of larger, well-curated datasets on model performance. GottBERT further extended this development as one of the first German RoBERTa models, trained on the German OSCAR corpus. Its results demonstrated the importance of data diversity but also noted that excessive data cleaning might reduce corpus variance and affect downstream performance. GeistBERT refines this lineage by improving data variance, optimizing pre-training strategies, and enhancing resource efficiency, making it a more robust model for German NLP.
@@ 3 Methodology
Compared to $\mathrm{GottBERT}$, GeistBERT was trained on a substantially larger corpus, totaling approximately 1.3TB of text data. The training data, listed in Table 1, was used in a shuffled version to enhance diversity and robustness during pre-training. GeistBERT was pre-trained using the same byte-level BPE tokenizer as $\mathrm{GottBERT}$, following the GPT-2 design with a vocabulary size of 52k. While the tokenizer architecture mirrors GPT-2, the vocabulary itself was trained from scratch on German text. Fairseq [23] was employed to compute the binary format for pre-training.
Using fairseq, we pre-trained the GeistBERT model on a highly variant corpus consisting of 1.3TB plain text data on 8 NVIDIA A40 GPUs. The model was trained with the RoBERTa base architecture for 100k update steps using a batch size of 8k, initializing the weights with ${}^{\mathrm{f}}\mathrm{GottBERT}_{base}$. We largely adhered to RoBERTa’s default training configuration [18], including dynamic masking, optimizer settings, and fixed sequence lengths (512 tokens). A 10k iteration warmup was applied, gradually increasing the learning rate to a peak of 0.0007, followed by a polynomial decay to zero. The pre-training process required approximately 8.3 days of computation time. In order to generate the Nyströmformer and Longformer version, we applied a post-hoc transformation to convert the resulting checkpoint for long-context use.
Table 1: Overview of the datasets used for training. The table lists the individual corpora, their sizes in gigabytes, their data sources, and whether they were deduplicated. The final corpus aggregates all listed datasets, resulting in approximately 1.3 TB of training data.
@@
We fine-tuned pre-trained BERT models using Huggingface [31] scripts, optimizing batch size and learning rate via grid search. NER and classification (CLS) tasks were trained for up to 30 epochs, while NLI tasks ran for up to 10 epochs using Fairseq-adapted hyperparameters. Each task was executed 24 times with varied hyperparameters, selecting the best checkpoint based on the highest $F_{1}$ score (accuracy for NLI). Performance was evaluated analogously to Scheible et al. [26] and compared with results from that study.
@@
We evaluated NLI on the German XNLI dataset [6], an extension of MultiNLI [30], with 122k training, 2490 development, and 5010 test examples per language. Performance was measured by accuracy.
@@
NER evaluation used the German CoNLL 2003 [29] and GermEval 2014 [3] datasets. CoNLL 2003 includes three entity classes, while GermEval 2014 provides fine-grained labels and nested markables. Both were evaluated using the $F_{1}$ score, with GermEval using an adapted metric accounting for label and span equality.
@@
We evaluated classification on GermEval 2018 [24] (German tweet sentiment analysis) and 10kGNAD [25] (German news categorization). GermEval 2018 followed Chan et al. [5] splits, while 10kGNAD used a predefined 90%-10% split with 10% for validation. Both tasks were assessed using the mean $F_{1}$ score.
@@ 4 Results
GeistBERT establishes a new state-of-the-art (SOTA) in German NLP, outperforming both base and large models in multiple tasks while maintaining efficiency and robustness. The results of all the downstream tasks are listed in Table 4.
@@
GeistBERTbase achieves 82.67% accuracy, surpassing all other base models. Its Longformer and Nyströmformer adaptations remain highly competitive, performing on par with or exceeding previous best-performing models.
Table 2: All the results of the experiments are shown in percent. They are all based on the test set and the best score out of 24 runs (selection based on validation set). While NLI is measured by accuracy, all the other metrics are $F_{1}$ measures. Best score in bold and second underlined, for large and base models respectively. For the $\mathrm{GottBERT}$ models, we report results on both the unfiltered and filtered corpora (marked by f). Each setup was evaluated per epoch, reporting the best (based on validation) and last checkpoints, with the latter marked by   . Values for non-GeistBERT models are taken from Scheible et al. [26].
@@
GeistBERTlongformer leads among base models, achieving 86.71% (CoNLL 03) and 88.45% (GermEval 2014), outperforming even some large models. GeistBERT’s strong results in both datasets validate its ability to capture complex entity relationships and maintain data diversity.
@@
GeistBERT and its variants dominate classification tasks, achieving 80.56% (GermEval 2018 coarse) and 90.89% (10kGNAD). Notably, GeistBERTnyströmformer surpasses base and many large models in fine-grained classification, showcasing its effectiveness in nuanced text understanding.
@@ 5.1 Principal Findings
The continued pre-training on GottBERT in combination with a broader, partly deduplicated German corpus (OSCAR23, OPUS, mC4, Wikipedia, OpenLegal) as well as the integration of WWM evidently leads to substantial improvements on various language modeling tasks. Effectively, this makes GeistBERT a novel SOTA model, surpassing even larger models in multiple German NLP tasks.
@@
Following GottBERT’s widespread adoption in NLP research [27, 4, 17], GeistBERT extends its usability in machine translation [33], NER [13, 12], and domain-specific applications. In contrast to the TPU-based training used for GottBERT, GeistBERT was pre-trained on GPUs, which allowed for more flexible data preprocessing, including sentence-aligned segmentation. This made it possible to preserve natural sentence structure during training, even when using fixed-length sequences. Nevertheless, hyperparameter tuning remains a crucial factor for achieving strong downstream performance [10]. WWM contributed to improved tokenization, aligning with previous findings [19, 5]. Certain Longformer-based tasks required multiple GPUs, even with gradient accumulation [22] steps to simulate larger batch sizes (48 and 64). In our experience, Longformer consumed significantly more VRAM compared to Nyströmformer and RoBERTa, which could be trained on a single RTX 3090 (24 GB VRAM) without gradient accumulation. Beyond memory consumption, inference speed and training time are important practical considerations. Table 7 provides a detailed comparison of training and evaluation time across GeistBERT models and tasks. As shown, Longformer consistently incurred higher computational costs—both during training and inference—and often required multiple GPUs, whereas Nyströmformer offered a more efficient trade-off between resource usage and performance. While deduplication was applied to specific subcorpora, the overall dataset was not filtered, meaning some redundant or lower-quality data may still be present. Prior work suggests that models benefit from increased corpus diversity [19], and GeistBERT’s use of many different corpora likely contributed to its robustness. Noteworthily, we chose to continue pre-training from $\mathrm{GottBERT}$ rather than training GeistBERT from scratch, as it is common practice with domain-specific adaptations [17, 16, 1, 14]. This decision allowed us to leverage pre-trained weights and rely on existing German-specific tokenization. Additionally, given the limited time available for this project, we prioritized the training and thorough evaluation of a continued pre-training approach as a more conservative and focused choice over training multiple models including one from scratch. El Boukkouri et al. [11] argue that training from scratch with a specialized vocabulary could offer more tailored embeddings. Their findings also indicate that continued pre-training on a domain-specific corpus yields comparable results in most cases. Although our strategy of relying on GottBERT’s pre-work led to strong performance across multiple downstream tasks, we acknowledge that a comparison to training from scratch on the same data and architecture would provide additional insight. We consider this a direction for future work. Another limitation of our study is the absence of long-context evaluation. This is primarily due to the lack of established benchmarks for long-context understanding in German. Creating such datasets is a complex and resource-intensive task, and is typically pursued as part of dedicated efforts within the NLP community. We consider this an important avenue for future work.
5.2 Limitations
@@
Due to efficiency constraints and limited computational resources, we did not train a large GeistBERT model. While our results demonstrate the strong performance of the base models, larger architectures could potentially achieve even better results.
5.3 Ethical Considerations
As with all large-scale NLP models, GeistBERT inherits biases from its training data, which could impact downstream applications. While deduplication reduces noise, it does not eliminate underlying biases. Additionally, privacy concerns arise from large web-based datasets, as models may retain sensitive information. Future research should explore privacy-preserving techniques to mitigate this risk.
Another consideration is the environmental cost of training large transformer models. While we optimized for efficiency, pre-training and evaluation remains computationally expensive, contributing to high energy consumption and carbon emissions.
As a RoBERTa-based model, GeistBERT does not generate free-form text but can still propagate biases present in its training data, potentially affecting downstream applications such as classification, sentiment analysis, or automated decision-making. Ensuring fairness, transparency, and responsible usage is essential, especially in sensitive domains such as legal, medical, or financial NLP.
6 Conclusion
In this work, we present GeistBERT, a German RoBERTa-based model trained on a diverse, partly deduplicated corpus, incorporating Whole Word Masking (WWM) and extended input capabilities. GeistBERT achieves state-of-the-art (SOTA) performance, outperforming even larger models in multiple tasks. Our findings highlight the importance of corpus diversity and advanced pre-training techniques in model effectiveness. All GeistBERT models are released on Huggingface, with the base model additionally in fairseq, under the MIT license, advancing German NLP.
Funding Declaration
This research did not receive any external funding.
Competing Interests
The authors have no competing interests to declare that are relevant to the content of this article.

### v4

@@ Abstract
Advances in transformer-based language models have highlighted the benefits of language-specific pre-training on high-quality corpora. In this context, German NLP stands to gain from updated architectures and modern datasets tailored to the linguistic characteristics of the German language. GeistBERT seeks to improve German language processing by incrementally training on a diverse corpus and optimizing model performance across various NLP tasks. We pre-trained GeistBERT using fairseq, following the RoBERTa base configuration with Whole Word Masking (WWM), and initialized from $\mathrm{GottBERT}$ weights. The model was trained on a 1.3 TB German corpus with dynamic masking and a fixed sequence length of 512 tokens. For evaluation, we fine-tuned the model on standard downstream tasks, including NER (CoNLL 2003, GermEval 2014), text classification (GermEval 2018 coarse/fine, 10kGNAD), and NLI (German XNLI), using $F_{1}$ score and accuracy as evaluation metrics. GeistBERT achieved strong results across all tasks, leading among base models and setting a new state-of-the-art (SOTA) in GermEval 2018 fine text classification. It also outperformed several larger models, particularly in classification benchmarks. To support research in German NLP, we release GeistBERT under the MIT license.
@@ 1 Introduction
The advancement of neural language modeling (LM) in natural language processing (NLP) has been driven by the development of contextual pre-trained word representations, particularly through transformer-based architectures. Models like Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019) have significantly impacted the field by providing robust, generalized representations that can be fine-tuned for specific downstream tasks, enhancing performance across various NLP applications. While much of the early work focused on English and multilingual models, it has become clear that single-language models, particularly those trained on large, high-quality corpora, can outperform their multilingual counterparts when applied to their target language.
Building on this understanding, the German NLP community has seen the introduction of models like GottBERT (Scheible et al., 2024), which leveraged the German portion of the OSCAR (Ortiz Suárez et al., 2020) corpus to create a high-performance RoBERTa-based (Liu et al., 2019) model tailored specifically for the German language. However, as the field evolves, so too must the approaches to model training. Recent developments in pre-training methodologies, such as Whole Word Masking (WWM) (Cui et al., 2021) and the availability of newer, more extensive corpora like OSCAR23 (Jansen et al., 2022), OPUS (Tiedemann, 2012), and mC4 (Xue et al., 2021), present opportunities to further refine and enhance German language models.
To fully leverage these developments for German NLP, we introduce GeistBERT, a German Enhanced Incremental Semantic Tuning BERT model. GeistBERT builds on the foundation laid by the best checkpoint of the filtered GottBERT model (i.e. ${}^{\mathrm{f}}\mathrm{GottBERT}_{base}$), incrementally pre-training on top of it using a combination of partly deduplicated modern German corpora, including OSCAR23, mC4 from CulturaX (Nguyen et al., 2023), Wikipedia and corpora from OPUS. By integrating WWM and harnessing the diverse, expansive nature of these contemporary corpora, GeistBERT aims to set a new benchmark for German language models, offering improved performance and broader applicability across a range of NLP tasks.
@@
  • We incrementally trained GeistBERT on top of ${}^{\mathrm{f}}\mathrm{GottBERT}_{base}$ using a combination of partly deduplicated modern German corpora (OSCAR23, OPUS, mC4), OpenLegal and Wikipedia.
  • We integrated WWM into the pre-training process to enhance the model’s ability to capture semantic relationships within the German language.
  • We provide GeistBERT as base model to the community, accessible under an open-source license for further usage.
GeistBERT represents a significant step forward in the development of German-specific transformer models, offering enhanced capabilities through modern training techniques and high-varying data.
@@ 2 Related Work
The rise of transformer-based models like BERT (Devlin et al., 2019) marked a major shift in NLP, enabling significant performance improvements. Originally introduced as an English model and later as a multilingual version (mBERT), BERT’s success led to monolingual adaptations tailored to specific languages. For German, models like GermanBERT [^1] https://www.deepset.ai/german-bert and dbmdz BERT [^2] https://huggingface.co/dbmdz/bert-base-german-uncased emerged, trained on datasets of 12GB–16GB, sourced from Wikipedia, news articles, and legal texts.
RoBERTa enhanced BERT by training on a larger 160GB corpus, optimizing the architecture, and removing next sentence prediction. This strategy was applied to other languages, resulting in models like CamemBERT (Martin et al., 2020) for French and RobBERT (Delobelle et al., 2020) for Dutch, highlighting the benefits of large, diverse training corpora and the use of language-specific vocabularies.
In German NLP, GBERT and GELECTRA (Chan et al., 2020) built on this progress by training on 145GB of the OSCAR corpus (Ortiz Suárez et al., 2020) and additional sources, surpassing earlier German BERT models. These advancements underscored the impact of larger, well-curated datasets on model performance. GottBERT further extended this development as one of the first German RoBERTa models, trained on the German OSCAR corpus. Its results demonstrated the importance of data diversity but also noted that excessive data cleaning might reduce corpus variance and affect downstream performance. GeistBERT refines this lineage by increasing data variance, optimizing pre-training strategies, and achieving strong performance without increasing model size, making it a robust and accessible model for German NLP.
@@ 3 Methodology
Compared to $\mathrm{GottBERT}$, GeistBERT was trained on a substantially larger corpus, totaling approximately 1.3TB of text data. Training data was shuffled to support uniform sampling and minimize order effects during pre-training. GeistBERT was pre-trained using the same byte-level BPE tokenizer as $\mathrm{GottBERT}$, following the GPT-2 design with a vocabulary size of 52k. While the tokenizer architecture mirrors GPT-2, the vocabulary itself was trained from scratch on German text. fairseq (Ott et al., 2019) was employed to compute the binary format for pre-training.
Table 1: Overview of datasets used for training. The table lists the individual corpora, their sizes in gigabytes, their data sources, and whether they were deduplicated. The final corpus aggregates all listed datasets, resulting in approximately 1.3 TB of training data.
@@
Using fairseq, we pre-trained the GeistBERT model on a highly variant corpus consisting of 1.3TB plain text data on 8 NVIDIA A40 GPUs. The model was trained with the RoBERTa base architecture for 100k update steps using a batch size of 8k, initializing the weights with ${}^{\mathrm{f}}\mathrm{GottBERT}_{base}$. We largely adhered to RoBERTa’s default training configuration (Liu et al., 2019), including dynamic masking, optimizer settings, and fixed sequence lengths (512 tokens). A 10k iteration warmup was applied, gradually increasing the learning rate to a peak of 0.0007, followed by a polynomial decay to zero. The pre-training process required approximately 8.3 days of computation time.
@@
We fine-tuned pre-trained BERT models using Huggingface (Wolf et al., 2019) scripts, optimizing batch size and learning rate via grid search. NER and classification (CLS) tasks were trained for up to 30 epochs, while NLI tasks ran for up to 10 epochs using fairseq-adapted hyperparameters. Each task was executed 24 times with varied hyperparameters, selecting the best checkpoint based on the highest $F_{1}$ score (accuracy for NLI). Performance was evaluated analogously to Scheible et al. (2024) and compared with results from that study.
@@
We evaluated NLI on the German XNLI dataset (Conneau et al., 2018), an extension of MultiNLI (Williams et al., 2018), with 122k training, 2490 development, and 5010 test examples per language. Performance was measured by accuracy.
@@
NER evaluation used the German CoNLL 2003 (Tjong Kim Sang and De Meulder, 2003) and GermEval 2014 (Benikova et al., 2014) datasets. CoNLL 2003 includes four entity types, while GermEval 2014 provides fine-grained categories and supports nested annotations. Both were evaluated using the $F_{1}$ score, with GermEval using an adapted metric accounting for label and span equality.
@@
We evaluated classification on GermEval 2018 (Risch et al., 2018) (German tweet sentiment analysis) and 10kGNAD (Schabus et al., 2017) (German news categorization). GermEval 2018 followed the data splits defined by Chan et al. (2020), while 10kGNAD used a predefined 90%-10% train–test split, with 10% of the training set further held out for validation. Both tasks were evaluated using the mean $F_{1}$ score.
@@ 4 Results
GeistBERT sets a new state-of-the-art among base models for German NLP, outperforming all comparable models and closely approaching large-scale model performance across tasks. It even achieves absolute SOTA in GermEval 2018 fine-grained classification (see Table 4).
@@
[…]
@@
GeistBERT achieves strong $F_{1}$ scores on both CoNLL 2003 (86.17%) and GermEval 2014 (88.47%), outperforming all other base models in our evaluation. It also surpasses all large-scale $\mathrm{GottBERT}$variants on GermEval 2014 and comes remarkably close on CoNLL 2003, with only a 0.11% gap to the lowest-scoring large variant. While top-performing large models such as GBERTlarge (87.19% on CoNLL) and XLM-Rlarge (88.83% on GermEval) remain ahead, GeistBERT narrows the performance gap significantly, demonstrating robust entity representation capabilities despite its compact size.
@@
GeistBERTbase achieves strong performance across all classification tasks, ranking first in GermEval 2018 fine-grained classification (66.42%), second in 10kGNAD (90.89%), and third in GermEval 2018 coarse (79.67%). It consistently outperforms all other base models and surpasses several large-scale models, particularly in the fine-grained setting. The results indicate that GeistBERT performs competitively across diverse classification benchmarks, despite being a base-sized model.
@@ 5 Discussion
The continued pre-training of GottBERT on a broader and partially deduplicated German corpus consisting of OSCAR23, OPUS, mC4, Wikipedia, and OpenLegal, together with the use of WWM, leads to clear improvements across multiple language modeling tasks. GeistBERT establishes a new state of the art among base models and achieves competitive results with larger models across multiple German NLP benchmarks.
@@
In contrast to the TPU-based training used for GottBERT, GeistBERT was pre-trained on GPUs, which allowed for more flexible data preprocessing, including sentence-aligned segmentation. This made it possible to preserve natural sentence structure during training, even when using fixed-length sequences. Nevertheless, hyperparameter tuning remains a crucial factor for achieving strong downstream performance (Dodge et al., 2020). WWM contributed to improved tokenization, aligning with previous findings (Martin et al., 2020; Chan et al., 2020). While deduplication was applied to specific subcorpora, the overall dataset was not filtered, meaning some redundant or lower-quality data may still be present. Prior work suggests that models benefit from increased corpus diversity (Martin et al., 2020), and GeistBERT’s use of many different corpora likely contributed to its robustness. Additionally, vocabulary size plays a role in performance (Toraman et al., 2023), though ours remains well-optimized. Noteworthily, we chose to continue pre-training from $\mathrm{GottBERT}$ rather than training GeistBERT from scratch, as it is common practice with domain-specific adaptations (Lentzen et al., 2022; Lee et al., 2019; Arefeva and Egger, 2022; Gururangan et al., 2020). This allowed us to reuse German-specific tokenization and pre-trained weights, and to focus on training and evaluating a single, well-defined setup within time constraints. While training from scratch with a custom vocabulary may yield more tailored embeddings (El Boukkouri et al., 2022), prior work suggests that continued pre-training often achieves comparable results. A direct comparison between continued pre-training and training from scratch on the same architecture and corpus remains an interesting avenue for future work. Following the broad adoption of GottBERT in German NLP (Scherrmann, 2023; Bressem et al., 2024; Lentzen et al., 2022; Xu et al., 2021; Frei et al., 2022; Frei and Kramer, 2023), we hope GeistBERT will be similarly received and applied.
6 Conclusion
In this work, we introduced GeistBERT, a German RoBERTa-based language model trained on a diverse and partially deduplicated corpus, incorporating WWM to enhance pre-training. GeistBERT achieves SOTA performance among base models and even outperforms several larger models across multiple tasks. These results underscore the importance of corpus diversity and WWM in improving downstream performance. GeistBERT is released under the MIT license on Huggingface and fairseq.
Limitations
@@ 6 Conclusion
Due to efficiency constraints and limited computational resources, we did not train a large version of GeistBERT, as pretraining based on GottBERT estimates would have required approximately 4.75 times more compute. While our results demonstrate the strong performance of the base model, larger architectures could potentially achieve even better results.
Ethical Considerations
Like all large-scale language models, GeistBERT may inherit biases from its training data, which can influence downstream tasks such as classification or decision-making. While deduplication reduces redundancy and noise, it does not remove deeper societal or representational biases. Furthermore, training on large web-based corpora raises privacy concerns, as models may inadvertently retain sensitive information. Responsible deployment is especially important in high-stakes domains like legal, medical, or financial NLP.
Despite optimizations for efficiency, pre-training and evaluating transformer models remain computationally demanding, contributing to energy use and carbon emissions. These environmental costs highlight the need for balancing model performance with sustainable development goals.